Mole fraction of water is 0.85 in sulphuric acid solution then molality of solution A) 9.8 B) 10.8 C) 10.58 D) 11.25

Grade:12th pass

4 Answers

IHate Ppts
6 years ago
The relationship between mole fraction and molality is given as
m=1000 X xB/(1-XB)X Molecular weoght of solvent
The tricky part in this question is observing which is solute and solvent.
Mole fraction of water is 0.85 which means it is not the solute but it is the SOLVENT.
We know that mole fraction of solute+mole fraction of solvent=1
Therefore, xB=0.15
Substitute it in given equation,
1000(0.15)/(0.85)(18)
Which is equal to (A) 9.8
 
Raghav Goyal
5 years ago
Let the mass of solute be xLet the mass of solvent be100-xThe moles in solute will be x/98 as molecular mass is 98The moles in solvent will be 100-x/18 as mass of water as solvent is 18Mole fracion-no of molesof solute/moles of solute +moles of solvent.X/98/100-x/18=85/100because of mole fracionAfter solving eq we will get apprpx answer
Raja singh
5 years ago
nw=0.85
So nacid=1-0.85=0.15
Now
m=1000×nacid/nw×molecular wt of water
m=1000×0.15/0.85×18
=9.8mol/kg
 
Yash Chourasiya
3 years ago
Dear Studnet

Given,
mole fraction of water in solution = 0.85
So moles of acid = 1 – 0.85 = 0.15
Therefore,

m = 1000×moles of acid /(nw×molecular wt. of water​)
m = 1000×0.15​/(0.85×18)
= 9.8 mol/kg

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya